OpenStudy (anonymous):

A 15 foot ladder is leaning against a house with its bottom 9 feet from the base of the house. What is the slope of the ladder?

OpenStudy (anonymous):

need help

OpenStudy (anonymous):

is the question multiple choice?

OpenStudy (anonymous):

well is it?

OpenStudy (anonymous):

a) 915 b) 24 c) 12 d) 43 e) 15

OpenStudy (anonymous):

that's the multiple choice

OpenStudy (anonymous):

thanks now I have something to work with

OpenStudy (anonymous):

well I got the answer C but then my instructor said I didn't answer the question asked

OpenStudy (anonymous):

so what is the question

OpenStudy (anonymous):

what is the slope of the ladder?

OpenStudy (anonymous):



OpenStudy (anonymous):

\[\cos(\theta)=\frac{9}{15}=\frac{3}{5}\] \[\theta=\cos^{-1}(\frac{3}{5})\] and a calculator

OpenStudy (anonymous):

ooooh i get it, they want "slope" as in the slope of the line, not the angle ok no worries

OpenStudy (anonymous):

the base of the triangle is 9, the hypotenuse is 15 so by pythagoras the height is 12

OpenStudy (anonymous):

yes im trying to figure out the slope. I got the answer is 12

OpenStudy (anonymous):

and \[\frac{12}{9}=\frac{4}{3}\]

OpenStudy (anonymous):

but the slope? I don't know

OpenStudy (anonymous):

slope is "rise over run" which in this case is \(\frac{12}{9}\)

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

hold on are these a) 915 b) 24 c) 12 d) 43 e) 15 your answer choices? because if so it makes no sense

OpenStudy (anonymous):

this is the multiple choice... a) 9/15 b) 24 c) 12 d) 4/3 e) 15

OpenStudy (anonymous):

so I guess the answer is d

OpenStudy (anonymous):

yes i would go with D

OpenStudy (anonymous):

lol ok thanks so much